xx IDScalarWaveElliptic

IDScalarWaveElliptic

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 charge  10 Scope: private01  REAL

11 Description: Charge of uniformly charged sphere

11 Range 10              01  Default: 1.0
11 0:*  10 Anything

11

11 outputxtmp10 Scope: private01  BOOLEAN

11 Description: Output intermediate variables

11           10              01  Default: no

11

11 radius  10 Scope: private01  REAL

11 Description: Radius of uniformly charged sphere

11 Range 10              01  Default: 1.0
11 0:*  10 Anything

11

11 solver  10 Scope: private01  STRING

11 Description: Name of elliptic solver

11 Range 10              01  Default: sor
11 .*  10 Anything

11

11 sorxmaxit 10 Scope: private01  INT

11 Description: Maximum number of iterations for SOR solve

11 Range 10              01  Default: 100
11 1:*  10 Positive

11

11 initialxdata10 Scope: shared from IDSCALARWAVE01  KEYWORD

11 Extends ranges:

11 charge  10 Uniformly charged sphere

11

7 Interfaces

General Implements: idscalarwaveelliptic

Inherits: grid wavetoy ellbase idscalarwave

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  ellcoeffs  10                    01  compact 11 0
11                    10  Mcoeff  01 description 11 Coefficients of the elliptic equation
11                    10  Ncoeff  01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  outvars  10                    01  compact 11 0
11                    10  temp  01 description 11 Temporary vars
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: EllBase.h

8 Schedule This section lists all the variables which are assigned storage by thorn CactusWave/IDScalarWaveElliptic. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE Scheduled Functions

WaveToyxInitialData x uniformcharge x initial data (uniform charge) for 3d wave equation x

11 x 11 Language: 11 fortran
11 x 11 Storage: 11 wavetoy::scalarevolve[3]
11 x 11 x  11 ellcoeffs
11 x 11 x  11 outvars
11 x 11 Type:  11 function